# CAMPOS II. The onset of protostellar disk substructures and planet formation

### v1

@@ 4.3 The protostellar disk substructures with and without central cavities
We found that some protostellar disks have a central cavity in Class I and flat-spectrum phase. To further investigate this, we divide the substructure in two more general groups. The cavity-type (C-type) and the centrally filled type (F-type). The cavity-type disk substructures exhibit large ($>$10 au) cleared central dust cavities. These C-type protostellar disk substructures display morphologies similar to the more evolved Class II/Class III transition disks (van der Marel et al., 2016b). The origin of the central dust cavity in transition disks can be attributed to a range of processes, like extensive particle growth (Tanaka et al., 2005; Dullemond & Dominik, 2005), photoevaporative mass-loss in winds (Clarke et al., 2001), dead zones (Regály et al., 2012) or embedded stellar or planetary companions (Ireland & Kraus, 2008). In particular, in systems with super-Jovian planets or stellar binaries that can clear the material in disks, deep gas gaps are expected inside cavities in the disk dust distribution (Price et al., 2018; Muley et al., 2019; van der Marel et al., 2021). The other type of disk substructure population is the centrally-filled type (F-type) disk substructure. In contrast to the C-type disk substructures, F-type disk substructures do not exhibit a central cavity.
Figure 6 shows that these 2 types of substructures as a function of evolutionary state. C-type disk substructures are found across all evolutionary stages from Class I to Class II. The detection of large cavities in younger Class I and Flat-spectrum sources cannot be explained by photoevaporative mass-loss in winds (Clarke et al., 2001). Instead, the central cavity in Class I can possibly be explained by radial drift (Birnstiel et al., 2010; Cridland et al., 2017). The dust particles will experience a headwind due to the gas drag, drifting towards the central protostar (Cridland et al., 2017). Another explanation for these central cavity is tidal interactions with companions. There is no evidence of binarity for any of the disks in Figure 6. We can rule out all binaries with orbital separations greater than 15 au, which corresponds to the observational linear resolution. Protostellar disks with a C-type disk substructures are good candidates to search for gap-opening protoplanets or close protobinaries.
@@ 5 Discussion
Figure 8: Left: The brightness temperature of the protostellar disks calculated from the full Planck equation as a function of the evolutionary stage (bolometric temperature). Right: The transition between disks with substructures and without substructures at $T_{\rm bol}\sim 200$ K is not due to the optical depth. We do not see a discontinuity for the brightness temperature to dust temperature ratio ($T_{\rm bright,Planck}\,/\,T_{\rm dust}$) at 200 K. For optically thick disks the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio should be close to 1. We adopted the dust temperature from Equation 1, and excluded the circumbinary disks.
@@
Since the discovery of rings and gaps in HL Tau protoplanetary disk (ALMA Partnership et al., 2015), many more evolved protoplanetary disks substructures (rings, gaps, spirals) have been detected (e.g., Long et al. 2018; Tsukagoshi et al. 2019; Pinilla et al. 2017; van der Plas et al. 2017; van der Marel et al. 2016a; Kudo et al. 2018; Pinilla et al. 2018; Guzmán et al. 2018; van Terwisga et al. 2018; Fedele et al. 2017; Leiendecker et al. 2022; Booth et al. 2024; Shi et al. 2024; Huang et al. 2024; Guerra-Alvarado et al. 2025). Guerra-Alvarado et al. (2025) recently observed over 73 Class II disks in the Lupus molecular cloud at a resolution of 4 au, and found that the substructure detection rate strongly depends on the effective resolution. For disks with sizes between 8 and 30 au ($2\leq\theta_{\rm D}/\theta_{\rm res}\leq 7.5$), the substructure detection rate is approximately 40 %, whereas it increases to $\sim$80for disks larger than 30 au ($\theta_{\rm D}/\theta_{\rm res}\geq 7.5$) (Guerra-Alvarado et al., 2025). Although many Class II disks are compact (Guerra-Alvarado et al., 2025), the majority of Class II protoplanetary disks observed at high linear (a few au) and effective resolution ($\theta_{\rm D}/\theta_{\rm res}\geq 7.5$) exhibit substructures (Andrews et al., 2018a; Huang et al., 2018; Bae et al., 2023; Guerra-Alvarado et al., 2025). The high substructure detection rate in well-resolved Class II disks indicates that substructures are already widespread by the Class II stage (Bae et al., 2023).
@@
The quest to search for disk substructures in the Class 0/I protostellar phase has generated a great interest in the disk community (Aso et al., 2021; Xu et al., 2023; Lee et al., 2023; Maureira et al., 2024b; Michel et al., 2023; Ohashi et al., 2023). In particular, the recent ALMA Large Program, eDisk, has made a significant contribution, surveying around 20 embedded Class 0/I protostellar disks and found three of the disks in their sample have substructures (L1489 IRS, Oph IRS 63 and IRAS 04169+2702) (Ohashi et al., 2023). Among these, substructures in Oph IRS 63 and L1489 IRS had been previously reported (Segura-Cox et al., 2020; Ohashi et al., 2022). Based on the prevalence of Class 0/I disks without substructures (detection rate 3 out of 20), Ohashi et al. (2023) concluded that substructures form rapidly near the Class I/II transition.
@@
Our analysis of 90 protostellar disks from CAMPOS, incorporating observations from eDISK, ODISEA, and DSHARP, shows that detection rates for Class I and Class II disks become comparable at $T_{\rm bol}\sim 200-400$ K. In this study, we accounted for potential biases, including variations in observational resolution and source inclination. Lower resolution and higher inclination can artificially reduce the visibility of disk substructures, disproportionately affecting the detection of younger disks in current high-resolution surveys. We also consider the possible systematic biases, such as the lack of extinction correction or Herschel data, on the bolometric temperature of each source. The onset of disk substructure at $T_{\rm bol}\sim 200-400$ K, corresponds to an approximate age of $\sim$0.2-0.4 Myr, which is much earlier than the age of $\sim 1\,$Myr – the expected age for when the transition between the Class I and Class II takes place (Evans et al., 2009; Dunham et al., 2015).
@@
It is important to note that the onset of disk substructures in this study could be an upper limit, and the actual onset of disk substructures may occur much earlier, within the Class 0 phase. We took a conservative approach and did not count sources that show small power variations as a function of distance in $uv-$space or small variations in the intensity radial profile but appear smooth in the image plane as disks with substructures (e.g., TMC-1A spiral [^6] We did not include TMC-1A as a substructure source because it was not identified as a disk with a spiral substructure in the eDisk survey (Ohashi et al., 2023), in contrast to Xu et al. (2023).: Aso et al. 2021; Xu et al. 2023, HH211 spiral [^7] HH211 is not included in Table 1, because it is an edge-on source.: Lee et al. (2023), Oph A SM1 gap [^8] Oph A SM1 is not included in Table 1 because it does not have a bolometric temperature measurement.: Maureira et al. (2024b), faint rings in Oph-emb-6, Oph-emb-9, and GSS 30 IRS 3999Ring-like structures are not visually identifiable by the naked eye in Oph-emb-6, Oph-emb-9, and GSS 30 IRS 3 (Michel et al., 2023; Hsieh et al., 2024).: Michel et al. (2023)). Our conservative approach minimizes the chances of false positives, but we may miss lower-contrast (shallower) substructures in younger disks with borderline detection. Furthermore, another possible reason for this onset being an upper limit is the choice to include all disks with $\theta_{D}/\theta_{\rm res}\geq 4$, since the detection rate is only $\sim 50\%$ for values between $3<\theta_{D}/\theta_{\rm res}\leq 10$ (Bae et al., 2023).
@@
In addition, optical depth can also affect the detection rate for the younger Class 0 or early Class I disks (e.g., Maureira et al., 2024a). To investigate this possibility further, we estimated the optical depth at 1.3 mm of our disks by computing the ratio between brightness temperature ($T_{\rm B,Planck}$) and the dust temperature. We estimated the brightness temperature of the disk from the average disk intensity using the full Planck equation. The estimated brightness temperature as a function of evolution stage ($T_{\rm bol}$) is plotted in the left panel of Figure 8. We then compared it with the dust temperature given by:
@@
For optically thick disks, the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio is expected to be close to 1. In Figure 8 right panel, we plot the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio as a function of evolutionary stage.
Although we do not find a sharp transition in optical depth between Class 0 and I, there is an indication of optical depth increasing from Class I back into the Class 0 stage. This supports that the zero detection rate in our large sample for $T_{\rm bol}\leq 163$ K might be affected by high optical depths at 1.3 mm. Further suggestive of this is the fact that the disks with detected substructures in our work seem to have, on average, less estimated optical depth than all the remaining disks with no substructures during the Class 0 and I stages. If the Class 0 protostellar disks are optically thick, then the onset of protostellar disk substructures should occur even earlier.
The rapid formation of disk substructures at early Class I phases suggests that protostars and protoplanets grow and evolve together. During the Class 0 and early Class I phases, the central protostar and protostellar disks are constantly fed by anisotropic infall via streamers (Pineda et al., 2020; Valdivia-Mena et al., 2022; Pineda et al., 2023). This anisotropic infall can perturb the protostellar disks, generating vortices and pressure bumps, resulting in substructure formation (Kuznetsova et al., 2022). It is possible that these earliest disk substructures are triggered by anisotropic infall (Kuznetsova et al., 2022; Zhao et al., 2025). Other mechanisms for substructure formation have also been proposed (see Bae et al. (2023) for a review). While we do not know the origin of these early disk substructures, these structures can possibly trigger subsequent giant planet formation (Lau et al., 2024). The early onset of protostellar disk substructures at $\sim$0.2 – 0.4 Myr is also consistent with recent population synthesis models, which show that rapid substructure formation within 0.4 Myr is needed to explain the observed spectral index distribution in Lupus protoplanetary disks (Delussu et al., 2024). Our findings of the early formation of disk substructures at early Class I phases confirm that star formation sets the initial conditions for planet formation in protostellar disks.

### v4

@@ 4.3 The protostellar disk substructures with and without central cavities
We found that some protostellar disks have a central cavity in Class I and flat-spectrum phase. To further investigate this, we divide the substructure into two more general groups. The cavity-type (C-type) and the centrally filled type (F-type). The cavity-type disk substructures exhibit large ($>$10 au) cleared central dust cavities. These C-type protostellar disk substructures display morphologies similar to the more evolved Class II/Class III transition disks (van der Marel et al., 2016b). The origin of the central dust cavity in transition disks can be attributed to a range of processes, like extensive particle growth (Tanaka et al., 2005; Dullemond & Dominik, 2005), photoevaporative mass-loss in winds (Clarke et al., 2001), dead zones (Regály et al., 2012) or embedded stellar or planetary companions (Ireland & Kraus, 2008). In particular, in systems with super-Jovian planets or stellar binaries that can clear the material in disks, deep gas gaps are expected inside cavities in the disk dust distribution (Price et al., 2018; Muley et al., 2019; van der Marel et al., 2021). The other type of disk substructure population is the centrally filled type (F-type) disk substructure. In contrast to the C-type disk substructures, F-type disk substructures do not exhibit a central cavity.
Figure 6 shows that these 2 types of substructures as a function of evolutionary state. C-type disk substructures are found across all evolutionary stages from Class I to Class II. The detection of large cavities in younger Class I and Flat-spectrum sources cannot be explained by photoevaporative mass-loss in winds (Clarke et al., 2001). Instead, the central cavity in Class I can possibly be explained by radial drift (Birnstiel et al., 2010; Cridland et al., 2017). In this scenario, dust particles experience a headwind due to the gas drag, drifting towards the central protostar (Cridland et al., 2017). Another explanation for these central cavities is tidal interactions with companions. The identification of binary systems in ALMA images relies on the presence of circumstellar disks around both components. While our ALMA data does not show evidence of binarity within the cavity of the C-type disks shown in Figure 6, we cannot exclude the presence of tight binary companions that either lack detectable dust disks or remain unresolved due to the linear resolution limit of 15 au. Protostellar disks with C-type disk substructures are good candidates to search for gap-opening protoplanets or close protobinaries.
@@
In Figure 8, we examine the prevalence of central cavities (C-type) in protostellar disks as a function of disk mass, disk radius, and evolutionary stage, as indicated by the bolometric temperature. For disks with substructures, each disk is assigned a value of 1 if it exhibits a central cavity (C-type) and 0 if it does not (F-type). We compute a moving average with a window size of 8, where the kernel slides through $T_{\rm bol}$, disk radius, and disk mass, and calculate the mean for every group of 8 disks. This mean represents the fraction of disks with a central cavity (C-type).
Figure 8 shows that, in general, protostellar disks with a central cavity are smaller and less massive than those without a central cavity. Classifying disks with substructures into those with and without large central cavities, we find both populations coexisting across evolutionary stages. The fraction of protostellar disks with a central cavity stays roughly constant at around $40\%$.
@@ 5 Discussion
5.1 Central cavities in protostellar disks: are disk substructures evolutionarily linked?
Cieza et al. (2021) proposed that the diversity of disk substructures is evolutionarily linked and can be explained by giant planet formation via core accretion, coupled with simple dust evolution. Building upon this framework, Orcajo et al. (2025) conducted numerical simulations of protoplanetary disks to further establish the unified evolutionary sequence of planet-induced substructures. The evolutionary scenario proposed by both studies outlines five distinct stages, defined primarily by morphological features.
Classifying disk morphologies serves as a critical first step in examining this evolutionary sequence, analogous to the Hubble tuning fork used for galaxy classification. However, unlike galaxy classification, which relies on easily distinguishable features such as spiral versus elliptical structures, differentiating between stages of annular substructures is considerably more challenging due to their subtler morphological distinctions.
To test whether or not protostellar disk substructures are evolutionarily linked, as argued by Cieza et al. (2021) and Orcajo et al. (2025), we adopt a binary classification scheme based on the presence (C-type) or absence (F-type) of a central cavity. This classification is advantageous due to its simplicity, robustness, and clarity. If disk substructures indeed evolve over time, as proposed, we expect F-type disks to dominate in the earlier stages, with an increasing prevalence of C-type disks at later evolutionary stages.
By classifying disks with substructures into those with (C-type) and without (F-type) central cavities, we find that both populations coexist across evolutionary stages (Figure 8). Our results suggest that there is no clear, uniform evolutionary sequence in disk substructures. If the evolution of the disk substructures follows the unified evolutionary sequence proposed by Cieza et al. (2021) and Orcajo et al. (2025), then the data suggest that disk substructures evolve very rapidly and thus can be present in all Class I/II stages and/or that they can be triggered at different times. Alternatively, rather than representing an evolutionary path, the observed diversity could be explained by multiple distinct populations of disk substructures, shaped by different formation mechanisms.
We found that disks with a central cavity are, in general, smaller in size and less massive than disks without a central cavity. Most multi-ring systems fall into the F-type category, and multi-ring systems with a large central cavity (occupying at least 25% of the disk diameter) are exceedingly rare. A few notable exceptions include ISO-Oph 2a (Hsieh et al., 2024; González-Ruilova et al., 2020) and HD 169142 (Pérez et al., 2019), though these are atypical cases. HD 169142, for instance, is significantly older ($\sim$10 Myr; Pohl et al. 2017) than the protostellar disks analyzed in this study (ages $\lesssim 1$ Myr). This disk also hosts an embedded Jupiter-mass protoplanet, which interacts with the disk and likely contributes to the formation of a central cavity (Law et al., 2023). The other exception, ISO-Oph 2a, when observed at higher angular resolution, reveals two concentric rings (González-Ruilova et al., 2020). Recent multi-frequency ALMA observations show that this disk exhibits azimuthal temperature variations likely caused by a fly-by interaction with a secondary companion (Casassus et al., 2023). The morphology of ISO-Oph 2a, shaped by dynamical interactions, differs significantly from that of the other multi-ring systems presented in Figure 7.
We speculate that the two populations of disk substructures are related to different planetary systems architectures, formed through different mechanisms, and will possibly evolve into different types of planetary systems. The recent discovery of the first protoplanetary system, PDS 70, revealed two giant planets located at 20 and 35 au from the host star (Keppler et al., 2018; Haffert et al., 2019). In contrast, our solar system is much more compact, with Jupiter at $\sim 5$ au and Saturn at $\sim 10$ au. Jupiter formed much closer to the Sun at $\sim 3$ au before entering a resonance with Saturn and eventually migrating to its current position, as suggested in the Grand Track Model (Walsh et al., 2011). The formation of Jupiter-like giant planets would create a strong pressure bump around the planet, blocking mm-sized dust grains and preventing the dust in the outer disk from replenishing the inner disk. Over time, this causes the inner disk to dissipate completely, forming a central cavity (Zhu et al., 2012; Pinilla et al., 2016). Thus, protostellar disks with C-type substructures might form compact planetary systems that resemble our Solar system.
On the other hand, the gaps and rings in the multiple-rings system imaged by ALMA (F-type) are located much further away, at 10s of au from the host star. The location of these gaps and rings contrasts with the distribution of giant planets’ orbits, which peaks at 3 au as measured by the California Legacy Survey (Fulton et al., 2021). These multiple-ring systems are also larger and more massive than the disks with a central cavity, possibly suggesting that they might be the precursors of extended planetary systems like PDS 70. Alternatively, simulations show that multiple gaps and rings can be generated by a single planet inside a low viscosity disk (Dong et al., 2017; Zhang et al., 2018) or explained by various instabilities (see Bae et al. 2023 for a review). We hypothesize that some of the F-type multiple-ring disks with shallower gaps originated from various instabilities or low-mass planets inside a low-viscosity disk. In contrast, disks with a central cavity (C-type) are more likely to be associated with giant planet formation as compared to disks without a central cavity.
5.2 The onset of protostellar disk substructures
@@
Figure 9: Left: The brightness temperature of the protostellar disks calculated from the full Planck equation as a function of the evolutionary stage (bolometric temperature). Right: The transition between disks with substructures and without substructures at $T_{\rm bol}\sim 200$ K is not due to the optical depth. We do not see a discontinuity for the brightness temperature to dust temperature ratio ($T_{\rm bright,Planck}\,/\,T_{\rm dust}$) at 200 K. For optically thick disks the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio should be close to 1 (black horizontal line). We adopted the dust temperature from Equation 1, and excluded the circumbinary disks.
@@
Since the discovery of rings and gaps in the HL Tau protoplanetary disk (ALMA Partnership et al., 2015), many more evolved protoplanetary disks substructures (rings, gaps, spirals) have been detected (e.g., Long et al. 2018; Tsukagoshi et al. 2019; Pinilla et al. 2017; van der Plas et al. 2017; van der Marel et al. 2016a; Kudo et al. 2018; Pinilla et al. 2018; Guzmán et al. 2018; van Terwisga et al. 2018; Fedele et al. 2017; Leiendecker et al. 2022; Booth et al. 2024; Shi et al. 2024; Huang et al. 2024; Guerra-Alvarado et al. 2025). Guerra-Alvarado et al. (2025) recently observed over 73 Class II disks in the Lupus molecular cloud at a resolution of 4 au, and found that the substructure detection rate strongly depends on the effective resolution. For disks with sizes between 8 and 30 au ($2\leq\theta_{\rm D}/\theta_{\rm res}\leq 7.5$), the substructure detection rate is approximately 40 %, whereas it increases to $\sim$80for disks larger than 30 au ($\theta_{\rm D}/\theta_{\rm res}\geq 7.5$) (Guerra-Alvarado et al., 2025). Although many Class II disks are compact (Guerra-Alvarado et al., 2025), the majority of Class II protoplanetary disks observed at high linear (a few au) and effective resolution ($\theta_{\rm D}/\theta_{\rm res}\geq 7.5$) exhibit substructures (Andrews et al., 2018a; Huang et al., 2018; Bae et al., 2023; Guerra-Alvarado et al., 2025). The high substructure detection rate in well-resolved Class II disks indicates that substructures are already widespread by the Class II stage (Bae et al., 2023).
@@
The quest to search for disk substructures in the Class 0/I protostellar phase has generated a great interest in the disk community (Aso et al., 2021; Xu et al., 2023; Lee et al., 2023; Maureira et al., 2024b; Michel et al., 2023; Ohashi et al., 2023). In particular, the recent ALMA Large Program, eDisk, has made a significant contribution, surveying around 20 embedded Class 0/I protostellar disks and found that three of the disks in their sample have substructures (L1489 IRS, Oph IRS 63 and IRAS 04169+2702) (Ohashi et al., 2023). Among these, substructures in Oph IRS 63 and L1489 IRS had been previously reported (Segura-Cox et al., 2020; Ohashi et al., 2022). Based on the prevalence of Class 0/I disks without substructures (detection rate 3 out of 20), Ohashi et al. (2023) concluded that substructures form rapidly near the Class I/II transition.
@@
Our analysis of 90 protostellar disks from CAMPOS, incorporating observations from eDISK, ODISEA, and DSHARP, shows that detection rates for Class I and Class II disks become comparable at $T_{\rm bol}\sim 200-400$ K. In this study, we accounted for potential biases, including variations in observational resolution and source inclination. Lower resolution and higher inclination can artificially reduce the visibility of disk substructures, disproportionately affecting the detection of younger disks in current high-resolution surveys. We also consider the possible systematic biases, such as the lack of extinction correction or Herschel data, on the bolometric temperature of each source. Our finding is consistent with the recent study of protostellar disks in the Ophiuchus molecular cloud by Shoshi et al. (2025), which found that disk substructure starts to appear when $T_{\rm bol}$ exceeds 200–300 K. The onset of disk substructure at $T_{\rm bol}\sim 200-400$ K, corresponds to an approximate age of $\sim$0.2-0.4 Myr, which is much earlier than the age of $\sim 1\,$Myr – the expected age for when the transition between the Class I and Class II takes place (Evans et al., 2009; Dunham et al., 2015).
@@
It is important to note that the onset of disk substructures in this study could be an upper limit, and the actual onset of disk substructures may occur much earlier, within the Class 0 phase. We took a conservative approach and did not count sources that show small power variations as a function of distance in $uv-$space or small variations in the intensity radial profile but appear smooth in the image plane as disks with substructures (e.g., TMC-1A spiral [^7] We did not include TMC-1A as a substructure source because it was not identified as a disk with a spiral substructure in the eDisk survey (Ohashi et al., 2023), in contrast to Xu et al. (2023).: Aso et al. 2021; Xu et al. 2023, HH211 spiral [^8] HH211 is not included in Table 1, because it is an edge-on source.: Lee et al. (2023), Oph A SM1 gap [^9] Oph A SM1 is not included in Table 1 because it does not have a bolometric temperature measurement.: Maureira et al. (2024b), faint rings in Oph-emb-6, Oph-emb-9, and GSS 30 IRS 3101010Ring-like structures are not visually identifiable by the naked eye in Oph-emb-6, Oph-emb-9, and GSS 30 IRS 3 (Michel et al., 2023; Hsieh et al., 2024).: Michel et al. (2023)). Our conservative approach minimizes the chances of false positives, but we may miss lower-contrast (shallower) substructures in younger disks with borderline detection. Furthermore, another possible reason for this onset being an upper limit is the choice to include all disks with $\theta_{D}/\theta_{\rm res}\geq 4$, since the detection rate is only $\sim 50\%$ for values between $3<\theta_{D}/\theta_{\rm res}\leq 10$ (Bae et al., 2023).
@@
In addition, optical depth can also affect the detection rate of substructures in younger Class 0 or early Class I disks (e.g., Maureira et al., 2024a). To investigate this possibility further, we estimated the optical depth at 1.3 mm of our disks by computing the ratio between brightness temperature ($T_{\rm B,Planck}$) and the dust temperature. We estimated the brightness temperature of the disk from the average disk intensity using the full Planck equation. The estimated brightness temperature as a function of evolution stage ($T_{\rm bol}$) is plotted in the left panel of Figure 9. We then compared it with the dust temperature given by:
@@
For optically thick disks, the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio is expected to be close to 1. In the right panel or Figure 9 we plot the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio as a function of evolutionary stage.
Although we do not find a sharp transition in optical depth between Class 0 and I, there is an indication of optical depth increasing from Class I back into the Class 0 stage. The median brightness temperature to dust temperature ratio ($T_{\rm bright,Planck}/T_{\rm dust}$) for Class 0 sources is 0.36, which is higher than that of Class I sources ($70\leq T_{\rm bol}\leq 650$ K), at 0.23, and Class II sources ($T_{\rm bol}\geq 650$ K), at 0.27. This supports that the zero detection rate in our large sample for $T_{\rm bol}\leq 163$ K might be affected by high optical depths at 1.3 mm. Further suggestive of this is the fact that the disks with detected substructures in our work seem to have, on average, less estimated optical depth than all the remaining disks with no substructures during the Class 0 and I stages. If the Class 0 protostellar disks are optically thick, then the onset of protostellar disk substructures likely occurs even earlier.
The rapid formation of disk substructures at early Class I phases suggests that protostars and protoplanets grow and evolve together. During the Class 0 and early Class I phases, the central protostar and protostellar disks are constantly fed by anisotropic infall via streamers (Pineda et al., 2020; Valdivia-Mena et al., 2022; Pineda et al., 2023). This anisotropic infall can perturb the protostellar disks, generating vortices and pressure bumps, resulting in substructure formation (Kuznetsova et al., 2022). These earliest disk substructures may be triggered by anisotropic infall (Kuznetsova et al., 2022; Zhao et al., 2025). Other mechanisms for substructure formation have also been proposed (see Bae et al. (2023) for a review). While we do not know the origin of these early disk substructures, these structures can possibly trigger subsequent giant planet formation (Lau et al., 2024). The early onset of protostellar disk substructures at $\sim$0.2 – 0.4 Myr is also consistent with recent population synthesis models, which show that rapid substructure formation within 0.4 Myr is needed to explain the observed spectral index distribution in Lupus protoplanetary disks (Delussu et al., 2024). Our findings of the early formation of disk substructures at early Class I phases confirm that star formation sets the initial conditions for planet formation in protostellar disks.